# R3-RAG: Learning Step-by-Step Reasoning and Retrieval for LLMs via Reinforcement Learning

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) integrates external knowledge with Large Language Models (LLMs) to enhance factual correctness and mitigate hallucination. However, dense retrievers often become the bottleneck of RAG systems due to their limited parameters compared to LLMs and their inability to perform step-by-step reasoning. While prompt-based iterative RAG attempts to address these limitations, it is constrained by human-designed workflows. To address these limitations, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving comprehensive external knowledge and leading to correct answers. R3-RAG is divided into two stages. We first use cold start to make the model learn the manner of iteratively interleaving reasoning and retrieval. Then we use reinforcement learning to further harness its ability to better explore the external retrieval environment. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcome reward, which judges whether the trajectory leads to a correct answer; 2) relevance-based document verification for process reward, encouraging the model to retrieve documents that are relevant to the user question, through which we can let the model learn how to iteratively reason and retrieve relevant documents to get the correct answer. Experimental results show that R3-RAG significantly outperforms baselines and can transfer well to different retrievers. We release R3-RAG at https://github.com/Yuan-Li-FNLP/R3-RAG.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a prevailing mechanism to address Large Language Models’ (LLMs) hallucination, as shown in Figure 1(a). It typically first retrieves relevant documents and then integrates them to supplement the LLM’s factual knowledge. In this way, RAG can supplement LLMs with external knowledge, mitigating hallucination and leading to better factuality.
@@
Figure 1: Comparison of different RAG approaches: (a) Vanilla RAG: The LLM uses the documents retrieved by the original question to generate the response; (b) Iterative RAG: The LLM interleaves between thinking and invoking the retriever in the fixed, human-designed workflow. (c) R3-RAG: Using reinforcement learning to let the LLM learn how to better reason and retrieve iteratively to get relevant documents and lead to the correct answer.
The response quality of RAG highly depends on the retrieval result  (Li et al., 2024; Yoran et al., 2024). However, the widely used retrievers have become the bottleneck of the RAG system and limited the overall performance. Specifically, dense retrievers usually have significantly fewer parameters and have not been scaled well to the size of LLMs. Additionally, dense retrievers directly generate vectors for input sequences and cannot conduct step-by-step thinking, which makes it challenging to tackle queries that require reasoning. RAG with iterative retrieval (see Figure 1(b)) is proposed to address these limitations. For example, Trivedi et al. (2023) propose to let the LLM interleave retrieval and Chain-of-Thought (CoT) to better retrieve relevant documents. Li et al. (2024) propose to let the LLM iteratively improve the retrieval result until it is sufficient to support answering the user question. However, these methods are limited by human-designed workflows and thus can not fully harness the LLM’s reasoning ability. Besides, LLMs are not trained on how to iteratively invoke retrievers to get comprehensive documents, making the human-designed workflow challenging to synergize the reasoning and retrieval.
To address these limitations, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving comprehensive external knowledge and leading to the correct answer. As shown in Figure 1(c), in R3-RAG, the model can explore better reasoning and retrieval trajectory by the feedback of outcome reward and process reward. Specifically, R3-RAG is divided into two stages. First, we leverage cold start to make the base model learn the manner of iterative reasoning and invoking the retriever. Since cold start makes it difficult for the model to fully explore the external retrieval environment and thus overshadows the model’s ability, we then use RL with outcome rewards to help the model learn how to better reason and retrieve external knowledge step-by-step. To encourage the model to retrieve documents that are relevant to the user question, we further propose relevance-based document verification as process rewards to measure each step’s relevance of reasoning and retrieval. In this way, the model cannot only learn how to obtain the correct answer through outcome rewards, but also reason and retrieve step-by-step through the fine-grained process rewards.
@@
  • To the best of our knowledge, R3-RAG is the first framework to let the LLM learn how to reason and retrieve step-by-step with outcome rewards and fine-grained process rewards.
  • We conduct comprehensive comparisons with extensive iterative RAG methods and results show that R3-RAG significantly outperforms them, thus indicating its effectiveness.
  • We have released our code and model to facilitate future research at https://github.com/Yuan-Li-FNLP/R3-RAG.
@@ 2 Related Work
The retrieval in RAG can be mainly divided into two categories: 1) Passive Retrieval Guu et al. (2020) implement the Retrieve-Read paradigm where retrievers fetch documents based on input queries before concatenating them with prompts for language model to generate responses, which is followed by many subsequent works  (Lewis et al., 2021; Gao et al., 2023; Li et al., 2023). However, dense retrievers cannot demonstrate reasoning step-by-step, and thus directly using the original question as the query makes it difficult to retrieve comprehensive external information. Iterative Retrieval is proposed to address this limitation, which enables interleaving between thinking and retrieving through decomposed queries. 2) Iterative Retrieval ReAct, proposed by Yao et al. (2023), synchronizes reasoning traces with retrieval actions, while Self-ask, proposed by Press et al. (2023), implements an autonomous question formulation mechanism during the reasoning process. ITER-RETGEN (Shao et al., 2023) introduces cyclic retrieval-generation patterns, FLARE (Jiang et al., 2023) incorporates adaptive retrieval when LLMs generate low-confidence tokens. IRCoT (Trivedi et al., 2023) strategically positions retrieval operations within chain-of-thought reasoning sequences. Compared with these iterative retrieval methods, which are highly limited by the human-designed workflow, R3-RAG can learn how to better explore the external retrieval environment through RL, which significantly addresses the limitation of the human-designed workflow.
@@
Recently we have seen the emergence of large reasoning models. OpenAI et al. (2024) introduces OpenAI-O1, which demonstrates remarkable step-by-step reasoning and reflection capabilities. Similarly, DeepSeek-AI et al. (2025) develops Deepseek-R1, which is trained by large-scale RL based on outcome rewards. Ma et al. (2025) introduces techniques that enable LLMs to verify their own reasoning steps, leading to more reliable problem-solving capabilities. However, these RL-based methods are highly limited by the base model’s knowledge  (Yue et al., 2025). In this paper, we explore making LLMs learn how to reason and retrieve external information step-by-step, overcoming knowledge limitations of models when addressing complex questions. Guan et al. (2025), Li et al. (2025), Song et al. (2025) and Zheng et al. (2025), concurrently to our work, propose to use RL to improve the iterative RAG. Our method differs in the critical reward design: While they only use the answer correctness as outcome reward, which cannot provide the fine-grained training signal, we additionally propose relevance-based document verification for process reward to better guide the model to iteratively retrieve relevant documents and thus obtain the comprehensive information for the question.
@@ 3 Method
RAG struggles to generate high-quality responses when retrievers fail to comprehensively retrieve the required knowledge. However, retrievers are usually weaker than LLMs, thus limiting the LLM’s remarkable ability. To address this limitation, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving the comprehensive external knowledge and generating the factual answer. Specifically, R3-RAG is divided into two stages, shown in Figure 2. First, we leverage cold start (DeepSeek-AI et al., 2025) to make the base model learn the manner of iterative reasoning and invoking retriever. Since cold start makes the model difficult to fully explore the external retrieval environment, we then use RL to further harness its ability to better search the relevant external knowledge. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcome reward, which judges whether the trajectory leads to correct answers by ground truth; 2) relevance-based document verification for proess reward, encouraging the model to retrieve documents that are relevant to the user question. In this section, we first introduce the trajectory of R3-RAG and then the cold start and RL training.
@@ 3.1 Trajectory Definition
In this section, we introduce the trajectory definition of R3-RAG. Given a multi-hop question-answering (QA) dataset $\mathcal{Q}=\{(q_{i},a_{i}^{*})\}_{i=1}^{n}$ where $q_{i}$ is a question and $a_{i}^{*}$ is its golden answer, we define a trajectory $T_{i}$ for each question $q_{i}$ as a sequence of reasoning and retrieval steps $T_{i}=\{s_{i}^{1},s_{i}^{2},\ldots,s_{i}^{|T_{i}|}\}$. Each step $s_{i}^{j}$ in trajectory $T_{i}$ is as follows:
@@
where the model iteratively reasons and invokes retrievers until it thinks the documents are enough for generating the answer at step $\lvert T_{i}\rvert$. $\boldsymbol{\xi_{i}^{j}}$ denotes the Reasoning process at step $j$, which analyzes all information, e.g., analysis and documents accumulated up to the step $j-1$ and plans the step $j$’s action; $\boldsymbol{\psi_{i}^{j}}$ is the Retrieval Query produced from this reasoning; $\boldsymbol{d_{i}^{j}}$ comprises the Documents returned by the retriever for that query; and $\boldsymbol{a_{i}}$ is the final Answer to the original question.
During the reasoning process $\boldsymbol{\xi_{i}^{j}}$, the model can step-by-step perform analysis of the information gathered in previous steps. If the existing documents are enough to support the question, the model will output the answer $\boldsymbol{a_{i}}$; otherwise, it determines what additional information is needed and invokes the retriever with a new query $\boldsymbol{\psi_{i}^{j}}$.
@@ 3.2 Cold Start
Since base models cannot yet interleave step-by-step reasoning with retrieval, we first employ a cold start stage for R3-RAG. We show the detailed cold start data synthesis in Algorithm 1. For each question, we build its response which follows the defined trajectory iteratively. In each iteration, we first let a strong LLM generate the reasoning step and corresponding action, divided into two cases: 1) the reasoning process and retrieval query, or 2) the reasoning process and final answer. If the LLM gives the final answer, which indicates the LLM thinks it has sufficient information to answer the question, we will finish the question’s response generation. After the response generation, each question has a corresponding trajectory of iterative reasoning and retrieving processes, accompanied by its predicted answer. Additionally, we perform rejective sampling (DeepSeek-AI et al., 2025) based on the question’s gold answer to filter out trajectories that lead to the wrong answer, which further improves the quality of cold-start data. Additional data generation details are available in Appendix A. By performing SFT with the cold-start dataset, we enable the model to learn the manner of iteratively reasoning and retrieval, which can help the model better sample positive trajectories and thus obtain higher RL training signal. Further training details are provided in Appendix B.
@@
1:  $\mathtt{MAX\_ITERATION}\leftarrow 5$$\mathcal{D}_{CS}\leftarrow\emptyset$
@@
3:     $T_{i},\mathtt{flag}\leftarrow[\,],\mathtt{False}$
@@
6:        if $\mathtt{{}^{\prime}answer^{\prime}}\in\mathtt{step}$ then
@@
9:        else if $\mathtt{{}^{\prime}query^{\prime}}\in\mathtt{step}$ then
10:           $\mathtt{step}[\mathtt{{}^{\prime}doc^{\prime}}]\leftarrow R(\mathtt{step}[\mathtt{{}^{\prime}query^{\prime}}])$

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) integrates external knowledge with Large Language Models (LLMs) to enhance factual correctness and mitigate hallucinations. However, dense retrievers often become the bottleneck of RAG systems due to their limited parameters compared to LLMs and their inability to perform step-by-step reasoning. While prompt-based iterative RAG attempts to address these limitations, it is constrained by human-designed workflows. To address these limitations, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step-by-step, thus retrieving comprehensive external knowledge and leading to correct answers. R3-RAG is divided into two stages. We first use cold start to make the model learn the manner of iteratively interleaving reasoning and retrieval. Then we use reinforcement learning (RL) to further harness its ability to better explore the external retrieval environment. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcome reward, which judges whether the trajectory leads to a correct answer; 2) relevance-based document verification for process reward, encouraging the model to retrieve documents that are relevant to the user question, through which we can let the model learn how to iteratively reason and retrieve relevant documents to get the correct answer. Experimental results show that R3-RAG significantly outperforms baselines and can transfer well to different retrievers.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a prevailing mechanism to address Large Language Models’ (LLMs) hallucinations, as shown in Figure 1(a). It typically first retrieves relevant documents and then integrates them to supplement the LLMs’ factual knowledge. In this way, RAG can supplement LLMs with external knowledge, mitigating hallucination and leading to better factuality.
@@
Figure 1: Comparison of different RAG approaches: (a) Vanilla RAG: the LLM uses the documents retrieved for the original question to generate the response; (b) Iterative RAG: the LLM interleaves thinking and invoking the retriever in a fixed, human-designed workflow; and (c) R3-RAG: uses reinforcement learning (RL) to enable the LLM to better reason and retrieve iteratively, to get relevant documents and produce the correct answer.
The response quality of RAG highly depends on the retrieval results (Li et al., 2024; Yoran et al., 2024). However, the widely used retrievers have become the bottleneck of the RAG system and have limited the overall performance. Specifically, dense retrievers usually have significantly fewer parameters and have not been scaled to the size of LLMs. Additionally, dense retrievers directly generate vectors for input sequences and cannot conduct step-by-step thinking, which makes it challenging to tackle queries that require reasoning. RAG with iterative retrieval (see Figure 1(b)) has been proposed to address these limitations. For example, Trivedi et al. (2023) propose that LLMs interleave retrieval and Chain-of-Thought (CoT) to better retrieve relevant documents. Li et al. (2024) propose that LLMs iteratively improve the retrieval results until they are sufficient to support answering users’ questions. However, these methods are limited by human-designed workflows and thus cannot fully harness LLMs’ reasoning ability. In addition, LLMs are not trained to iteratively invoke retrievers to obtain comprehensive documents, making it challenging for the human-designed workflow to synergize reasoning and retrieval.
To address these limitations, we propose R3-RAG, which uses Reinforcement learning (RL) to make the LLM learn how to Reason and Retrieve step-by-step, thus retrieving comprehensive external knowledge and leading to the correct answer. As shown in Figure 1(c), in R3-RAG, the model can explore better reasoning and retrieval trajectories using feedback from the outcome reward and the process reward. Specifically, R3-RAG is divided into two stages. First, we use a cold start to make the base model learn the manner of iterative reasoning and invoking the retriever. Since a cold start makes it difficult for the model to fully explore the external retrieval environment and thus overshadows the model’s ability, we then use RL with the outcome reward to help the model learn how to better reason and retrieve external knowledge step-by-step. To encourage the model to retrieve documents that are relevant to the user’s question, we further propose relevance-based document verification as a process reward to measure the relevance of each step’s reasoning and retrieval. In this way, the model not only learns how to obtain the correct answer through the outcome reward, but also learns to reason and retrieve step-by-step through the fine-grained process reward.
@@
  • To the best of our knowledge, R3-RAG is the first framework that enables LLMs to perform step-by-step reasoning and retrieval, guided by the outcome reward and the fine-grained process reward.
  • We conduct comprehensive comparisons with extensive iterative RAG methods, and the results show that R3-RAG significantly outperforms them, indicating its effectiveness.
  • We will release our code and model to facilitate future research.
@@ 2 Related Work
The retrieval in RAG can be mainly divided into two categories: 1) Passive Retrieval: Guu et al. (2020) implement the Retrieve-Read paradigm, where retrievers fetch documents based on input queries before concatenating them with prompts for LLMs to generate responses, a setup followed by many subsequent works (Lewis et al., 2021; Gao et al., 2023; Li et al., 2023). However, dense retrievers cannot perform step-by-step reasoning, and thus, directly using the original question as the query makes it difficult to retrieve comprehensive external information. Iterative retrieval has been proposed to address this limitation; it enables interleaving reasoning and retrieval through decomposed queries. 2) Iterative Retrieval: ReAct, proposed by Yao et al. (2023), synchronizes reasoning traces with retrieval actions, while Self-Ask, proposed by Press et al. (2023), implements an autonomous question formulation mechanism during the reasoning process. ITER-RETGEN (Shao et al., 2023) introduces cyclic retrieval-generation patterns; FLARE (Jiang et al., 2023) incorporates adaptive retrieval when LLMs generate low-confidence tokens; and IRCoT (Trivedi et al., 2023) strategically positions retrieval operations within CoT reasoning sequences. Compared with these iterative retrieval methods, which are limited by human-designed workflows, R3-RAG learns to better explore the external retrieval environment via RL, thereby alleviating this limitation.
@@
Recently, we have seen the emergence of large reasoning models. OpenAI et al. (2024) introduces OpenAI-O1, which demonstrates strong step-by-step reasoning and reflection capabilities. Similarly, DeepSeek-AI et al. (2025) develops Deepseek-R1, trained with large-scale RL based on outcome rewards. Ma et al. (2025) introduces techniques that enable LLMs to verify their own reasoning steps, leading to more reliable problem-solving capabilities. However, these RL-based methods are highly limited by the base model’s knowledge (Yue et al., 2025). In this paper, we explore training LLMs to reason and retrieve external information step-by-step, overcoming the base model’s knowledge limitations when addressing complex questions. Guan et al. (2025), Li et al. (2025), Song et al. (2025), and Zheng et al. (2025), concurrent with our work, propose using RL to improve iterative RAG. Our method differs in the critical reward design: while they use only answer correctness as an outcome reward, which cannot provide a fine-grained training signal, we additionally propose relevance-based document verification as a process reward to better guide the model to iteratively retrieve relevant documents and thus obtain comprehensive information for the question.
@@ 3 Method
RAG struggles to generate high-quality responses when retrievers fail to comprehensively retrieve the required knowledge. However, retrievers are usually weaker than LLMs, thus limiting the LLM’s ability. To address this limitation, we propose R3-RAG, which uses Reinforcement learning to enable the LLM to learn to Reason and Retrieve step-by-step, thereby retrieving comprehensive external knowledge and generating a factual answer. Specifically, R3-RAG is divided into two stages, as shown in Figure 2. First, we leverage cold start (DeepSeek-AI et al., 2025) to teach the base model the manner of iterative reasoning and invoking the retriever. Since a cold start makes it difficult for the model to fully explore the external retrieval environment, we then use RL to further harness its ability to better search for relevant external knowledge. Specifically, we propose two rewards for R3-RAG: 1) an answer-correctness outcome reward, which judges whether the trajectory leads to a correct answer against the ground truth; and 2) relevance-based document verification as a process reward, encouraging the model to retrieve documents that are relevant to the user’s question. In this section, we first introduce the R3-RAG trajectory and then the cold start and RL training.
@@ 3.1 Trajectory Definition
In this section, we define the R3-RAG trajectory. Given a multi-hop question-answering (QA) dataset $\mathcal{Q}=\{(q_{i},a_{i}^{*})\}_{i=1}^{n}$, where $q_{i}$ is a question and $a_{i}^{*}$ is its ground-truth answer, we define a trajectory $T_{i}$ for each question $q_{i}$ as a sequence of reasoning and retrieval steps $T_{i}=\{s_{i}^{1},s_{i}^{2},\ldots,s_{i}^{|T_{i}|}\}$. Each step $s_{i}^{j}$ in trajectory $T_{i}$ is as follows:
@@
where the model iteratively reasons and invokes the retriever until it deems the documents sufficient to generate the answer at step $\lvert T_{i}\rvert$. $\boldsymbol{\xi_{i}^{j}}$ denotes the Reasoning Process at step $j$, which analyzes all information (e.g., analysis and documents) accumulated up to step j1j-1, and plans the action for step $j$; $\boldsymbol{\psi_{i}^{j}}$ is the Retrieval Query produced by this reasoning; $\boldsymbol{d_{i}^{j}}$ comprises the Documents returned by the retriever for that query; and $\boldsymbol{a_{i}}$ is the final Answer to the original question.
During the reasoning process $\boldsymbol{\xi_{i}^{j}}$, the model can perform step-by-step analysis of the information gathered in previous steps. If the existing documents are sufficient to answer the question, the model outputs the answer $\boldsymbol{a_{i}}$; otherwise, it determines what additional information is needed and invokes the retriever with a new query $\boldsymbol{\psi_{i}^{j}}$.
@@ 3.2 Cold Start
Base models do not natively interleave step-by-step reasoning with retrieval. We therefore begin with a cold-start stage that synthesizes supervised trajectories to teach the desired interleaved behavior.
@@
[…]
@@
3:  $T_{i}\leftarrow[\,]$
@@
6:   if $\texttt{answer}\in\mathtt{step}$ then
@@
9:   else if $\texttt{query}\in\mathtt{step}$ then
10:    $\mathtt{step}[\texttt{doc}]\leftarrow R(\mathtt{step}[\texttt{query}])$
@@
3.2.1 Cold-Start Data Generation
Algorithm 1 outlines the procedure. For each question, we iteratively construct a trajectory whose steps follow a fixed schema. At each iteration, a powerful LLM is prompted to produce 1) a reasoning step with a retrieval query, or 2) a reasoning step with a final answer. For the first step, the template in Figure 6 prompts the model to analyze the problem, outline a solution path, and decide whether to retrieve or answer directly. For subsequent steps, the template in Figure 7 asks the model to consolidate prior reasoning and retrieved documents, assess the sufficiency of evidence, identify remaining knowledge gaps, and decide whether further retrieval is necessary or a final answer can be produced. If a final answer is produced, trajectory construction terminates. The resulting instance pairs each question with an interleaved reasoning–retrieval trajectory.
To improve data quality, we vary the sampling temperature and generate multiple trajectories per question, then perform rejection sampling against the ground-truth answer, discarding trajectories whose final answer is incorrect or whose steps violate the required format. The remaining trajectories constitute the cold-start dataset $\mathcal{D}_{\mathrm{CS}}$.
3.2.2 Cold-Start Training
To use $\mathcal{D}_{CS}$ for training in the cold-start stage—i.e., supervised fine-tuning (SFT)—we design Algorithm 2 to convert $\mathcal{D}_{CS}$ into an SFT-ready dataset $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m}$, where $m=\sum_{i=1}^{|\mathcal{Q}_{CS}|}|T_{i}|$. For each tuple $(q_{i},a_{i}^{*},T_{i})\in\mathcal{D}_{CS}$, we generate $|T_{i}|$ input–output pairs. Consider a step $s_{i}^{j}$ in trajectory $T_{i}$. We construct the input by concatenating the question $q_{i}$ with all steps that precede $s_{i}^{j}$, serialized into our unified format, including the reasoning (analysis $\xi_{i}^{j^{\prime}}$ and retrieval query $\psi_{i}^{j^{\prime}}$) and the retrieved documents $d_{i}^{j^{\prime}}$ for all $j^{\prime}<j$. We then construct the output as the reasoning content of the current step $s_{i}^{j}$ serialized in the same format. Iterating over all steps in $T_{i}$ yields $|T_{i}|$ training instances, which we add to $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}$.
Algorithm 2 Training Data Generation for SFT
0: Trajectory-augmented dataset $\mathcal{D}_{CS}$
0: SFT training dataset $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}$
[…]
2:for each $(q_{i},a_{i}^{*},T_{i})\in\mathcal{D}_{CS}$ do
3:  $\mathtt{context}\leftarrow q_{i}$
4:  $\mathtt{history}\leftarrow\emptyset$
5:  for $j=1$ to $|T_{i}|$ do
6:   $\mathtt{input}\leftarrow\mathtt{Concat}(\mathtt{context},\mathtt{history})$
7:   $s_{i}^{j}\leftarrow T_{i}[j]$
8:   if $j<|T_{i}|$ then
9:    $(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})\leftarrow s_{i}^{j}$
10:    $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,\psi_{i}^{j})$
11:    $\mathtt{step}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})$
12:    $\mathtt{history}\leftarrow\mathtt{Concat}(\mathtt{history},\mathtt{step})$
13:   else
14:    $(\xi_{i}^{j},\,a_{i})\leftarrow s_{i}^{j}$
15:    $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,a_{i})$
16:   end if
17:   $\mathtt{pair}\leftarrow(\mathtt{input},\mathtt{output})$
18:   $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}\leftarrow\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}\cup\{\mathtt{pair}\}$
19:  end for
20:end for
21:return $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}$
Based on $\mathcal{D}_{\mathrm{CS}}^{\mathrm{train}}$, we perform SFT with the maximum likelihood estimation objective to obtain the R3-RAG-CS model. This initialization enables the model to iteratively reason and retrieve, which can help the model better sample positive trajectories and thus yield stronger RL training signals.